. A square metal plate with a length of 0.20 m and with a charge of 9.7 x 10 -7 C.

Find the electric field of the square metal plate.


1
Expert's answer
2021-04-06T18:25:35-0400

Answer

Charge density of plate

"\sigma =\frac{Q}{A}"


"\sigma =\frac{9.7\times10^{-7}}{0.20^2}=24.25\mu C/m^2"



Electric fiele

"E=\frac{\sigma}{2\epsilon_0}=\frac{24.25\times 10^{-6}}{2(8.85\times10^{-12})}=1.37\times10^{6}N/C"